The slope of a line parallel to -x + y = 12 equals_______________,: The slope of a line parallel to -x + y = 12 equals_______________,
Answer by vleith(1238)

Put the line into the form y = mx + b
-x + y = 12
y = x + 12
Now it is in the form y = mx + b.
So what is m?
m=1 . The slope of this line is 1.
The slope of any line parallel to this line is also 1, since parallel lines have the same slope.